OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity 4 times x all over y all to the second power. 4 times x to the second power all over y to the second power 8 times x to the second power all over y to the second power 16 times x to the second power all over y to the second power 16 times x to the second power all over y

jimthompson5910 (jim_thompson5910):

can you draw out the original problem?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

You basically square each piece like this \[\large \left(\frac{4x}{y}\right)^2\] \[\large \frac{(4x)^2}{y^2}\] \[\large \frac{4^2*x^2}{y^2}\] \[\large \frac{16x^2}{y^2}\]

OpenStudy (anonymous):

thank you! can you help me with one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Choose the correct simplification of the expression 9 over h to the power of negative 3. h to the 3rd power over 9 9 over h to the 3rd power 9h3 This expression is already simplified

OpenStudy (anonymous):

 9 over h ^-3